It's a math problem what is the smallest number 1/5 3/10 4/15 3/16
Answer by Razzak Ahmed(10)

change them all to decimals first 1/5=.20, 3/10=.30, 4/15=.26, and 3/16=.18. From seeing you can tell that 3/16 is the smallest number.